# Taming the Tri-Space Tension: ARC-Guided Hallucination Modeling and Control for Text-to-Image Generation

### v1

@@ 5.2 Alignment Risk Code: Quantifying Real-time Tension Dynamics
We define the Alignment Risk Code (ARC) as a vector of instantaneous cognitive tensions at step $t$:
@@
The ARC vector provides a real-time, physically grounded representation of the internal alignment state of the model. From this vector, we derive interpretable measures:
  • Tension magnitude: $\|\vec{\tau}(p,t)\|_{2}$, indicating the overall alignment stress.
  • Tension imbalance: $\mathrm{Var}(\vec{\tau}(p,t))$, measuring the anisotropy of tension distribution.
  • Tension skew: $\mathrm{Softmax}(\vec{\tau})$, yielding a probability-like attribution of hallucination directionality.
Tracking $\vec{\tau}$ across diffusion steps reveals dynamic shifts in alignment priorities and exposes emerging risks of hallucination. As illustrated in Figure 4, certain prompts induce consistent dominance in specific $\tau_{i}$ components, leading to failure modes that are not random but axis-dependent.
The ARC formulation thus transforms latent alignment pressures into a structured, actionable representation. In the next section, we leverage this vector to develop a control mechanism that dynamically mitigates hallucinations by counteracting emerging tensions in real time, without requiring retraining of the base diffusion model.
@@ 6 TensionModulator: ARC-Guided Hallucination Control Mechanism
While the Alignment Risk Code (ARC) provides a real-time representation of cognitive tension during generation, it also enables a closed-loop correction framework. We introduce TensionModulator (TM-ARC), a lightweight, modular controller that applies dynamic perturbations to the generative trajectory in response to instantaneous tension signals. TM-ARC transforms ARC from a descriptive diagnostic vector into an actionable feedback mechanism, actively preventing directional drift toward the ideal manifold $\mathcal{M}_{\text{ideal}}$.
@@ 6.2 Modular Decomposition by Tension Axes
TM-ARC decomposes into three orthogonal sub-controllers, each aligned with a distinct cognitive alignment dimension:
  • SC-Gate ($F_{SC}$): Reactivates attention on prompt-critical entities when semantic tension $\tau_{SC}$ intensifies, counteracting semantic drift through modulation of decoder focus.
  • SA-Tuner ($F_{SA}$): Refines latent spatial encodings when structural tension $\tau_{SA}$ rises, preserving object placement and spatial logic via positional re-weighting.
  • KG-Augment ($F_{KG}$): Injects auxiliary factual priors or semantic anchors when knowledge tension $\tau_{KG}$ surges, reinforcing commonsense consistency. It operates in two complementary modes:
    • Static injection: KG-related prompt embeddings are prepended to the text encoder input to provide prior knowledge.
    • Dynamic modulation: Cross-attention layers are reweighted based on $\tau_{KG}$ to emphasize or suppress contextually relevant facts during generation.
Each submodule operates as a soft perturbation function over the latent state $z_{t}$, guided solely by the corresponding ARC component. The modular structure ensures that each intervention remains interpretable and independently tunable.
@@ 6.3 ARC-Driven Feedback Dynamics
  • $\lambda(\vec{\tau})=\sigma(\|\vec{\tau}\|_{2})$ is a tension-adaptive scaling function that increases modulation strength with overall tension magnitude;
  • $F_{i}(z_{t},\tau_{i})$ denotes the $i$-th correction operator, each modulated by its corresponding tension signal;
  • Each $F_{i}$ is differentiable and temporally localized, ensuring responsiveness without introducing global disruption.
This formulation allows TM-ARC to function analogously to a physical damping system: minimal corrections under low tension, and strong directional feedback when misalignment intensifies.
@@ 6.4 Integration and Generalization
TM-ARC is designed as a plug-and-play augmentation for pretrained diffusion models. It operates entirely in the latent space and requires no additional supervision, training, or architecture modification. The modular design ensures compatibility across model backbones and preserves zero-shot generalizability. Because its intervention is conditioned solely on ARC: a model-internal, prompt-dependent signal, TM-ARC generalizes across unseen prompts and domains, as long as alignment tension can be monitored.
By continuously sensing the evolving semantic, structural, and knowledge alignment tensions during generation, TM-ARC actively intervenes to guide the generative path back toward a cognitively balanced state. Thus, TM-ARC enables the first tension-aware feedback mechanism for hallucination mitigation in T2I diffusion, grounded in cognitive alignment dynamics rather than post hoc filtering or static constraints.
@@ 7 Experiments
We conduct comprehensive experiments to evaluate both the representational capacity of the proposed Alignment Risk Code (ARC) and the hallucination mitigation performance of the TensionModulator (TM-ARC) controller. Our analysis focuses on four key questions: (i) Can ARC capture structured hallucination patterns in a low-dimensional and interpretable manner? (ii) Does modeling multi-axis tension provide tangible benefits over lower-dimensional baselines? (iii) Can ARC enable prompt-aware, real-time controllability during generation? (iv) Does the TM-ARC module scale to diverse diffusion backbones and benchmarks while improving semantic alignment and perceptual quality? The following subsections progressively address these questions through quantitative metrics, visualization analyses, and large-scale benchmark studies.
7.1 3D ARC Yields the Most Expressive and Separable Representation
We first evaluate ARC’s ability to capture hallucination-specific alignment tensions in an interpretable and discriminative manner. As shown in Table 2, ARC’s full 3D formulation, spanning semantic coherence, structural alignment, and knowledge grounding, achieves a classification accuracy of 91.6% and a silhouette score of 0.63. In contrast, collapsing ARC to 2D or 1D results in significant performance drops, indicating that hallucination patterns cannot be adequately captured by any subset of axes alone. This supports our hypothesis that hallucinations are inherently multi-axial, requiring complete cognitive tension modeling.
7.2 Axis-Dominant Tension Explains Hallucination Failures and Guides Targeted Correction
Building on ARC’s representational capacity, we visualize hallucination decomposition across real prompts in Figure 6. Each example pair: faithful vs. hallucinated image, is accompanied by its ARC radar plot. Across diverse failure modes, hallucinations consistently align with specific tension spikes: semantic drift (e.g., ”puppy” $\rightarrow$ ”cat”) manifests as elevated $\tau_{\text{SC}}$; structural mismatch (e.g., object placement errors) correlates with $\tau_{\text{SA}}$; factual grounding issues (e.g., hallucinated ”AI conference”) are reflected in $\tau_{\text{KG}}$. Notably, our controller reduces the dominant spike while preserving balanced axes, showing that ARC is not only descriptive but also actionable.
@@ 7.3 Prompt-Level ARC Modulation Substantially Boosts Faithfulness
We next validate ARC’s role in enabling controllable generation. Table 3 shows that after TM-ARC modulation, ’Flying Elephant’ and ’Red Triangle’, both high-tension prompts, achieve +14.3 and +9.7 gains in faithfulness, respectively. The ARC vectors before and after intervention illustrate how the controller selectively attenuates risk-prone components. This demonstrates that ARC serves as a lightweight alignment monitor, enabling real-time axis-specific modulation during generation without retraining.
@@ 7.4 ARC Generalizes Across Backbones and Outperforms Strong Baselines
Table 2: Dimensionality ablation results on ARC representation.
ARC Dimension Classification Accuracy Silhouette Score
1D (best axis only) 78.2% 0.42
2D (any two axes) 86.7% 0.52
3D (full ARC) 91.6% 0.63
Table 3: ARC modulation and controllability performance.
Prompt ARC (Before) ARC (After) Faith $\uparrow$
Flying Elephant [0.92, 0.31, 0.41] [0.48, 0.30, 0.35] +14.3
Red Triangle [0.33, 0.88, 0.27] [0.29, 0.43, 0.24] +9.7
Table 4: Benchmark results on DrawBench and Pick-a-Pic. Best results are bolded; second-best are underlined. Metrics: CLIPScore (↑), PickScore (↑), ImageReward (IR↑), FID (↓).
Method CLIP Score↑ Pick Score↑ IR↑ FID↓ CLIP Score↑ Pick Score↑ IR↑ FID↓
DrawBench – SDXL DrawBench – SD1.5
Vanilla 28.3 20.9 0.77 47.5 28.7 21.7 0.86 20.0
Prompt-to-Prompt 27.4 20.8 0.70 27.0 27.5 20.4 0.59 54.0
Attend-and-Excite 28.4 20.6 0.68 47.0 29.9 22.5 0.58 26.5
Zigzag Diff 28.9 21.1 0.89 52.0 27.4 22.6 0.79 42.5
ARC (Ours) 29.3 21.6 0.85 20.0 28.7 21.1 0.89 20.0
DrawBench – PixArt-sigma DrawBench – Hunyuan-DiT
Vanilla 29.2 20.9 0.63 39.0 29.1 21.7 0.80 57.0
Prompt-to-Prompt 27.6 20.6 0.77 58.5 28.5 20.0 0.83 33.0
Attend-and-Excite 27.3 21.5 0.71 30.5 29.5 20.0 0.83 54.0
Zigzag Diff 29.1 21.6 0.75 24.6 28.9 20.2 0.83 49.5
ARC (Ours) 29.3 21.5 0.79 19.8 29.4 22.6 0.81 18.2
Pick-a-Pic – SDXL Pick-a-Pic – SD1.5
Vanilla 29.4 21.0 0.71 41.5 28.8 21.9 0.72 68.0
Prompt-to-Prompt 27.4 21.6 0.70 34.5 29.9 21.5 0.58 47.0
Attend-and-Excite 28.3 22.6 0.76 36.5 28.8 22.6 0.72 33.5
Zigzag Diff 29.1 21.8 0.86 20.5 28.2 22.5 0.79 29.5
ARC (Ours) 29.4 22.0 0.90 21.5 29.1 22.4 0.83 19.8
Pick-a-Pic – PixArt-sigma Pick-a-Pic – Hunyuan-DiT
Vanilla 27.9 22.5 0.68 29.0 27.8 22.9 0.78 27.0
Prompt-to-Prompt 29.8 21.5 0.57 32.5 27.5 21.7 0.82 35.5
Attend-and-Excite 29.8 20.8 0.59 50.0 28.5 22.9 0.71 69.0
Zigzag Diff 28.6 22.1 0.81 24.6 27.2 21.5 0.69 26.0
ARC (Ours) 29.7 22.5 0.83 17.7 29.1 22.7 0.83 18.8

### v3

@@ 3.1 Alignment Risk Code: Quantifying Real-time Tension Dynamics
We define the ARC as a vector of instantaneous cognitive tensions at step $t$:
@@
The ARC vector provides a real-time, physically grounded representation of the internal alignment state of the model. From this vector, we derive interpretable measures:(1) Tension magnitude: $\|\vec{\tau}(p,t)\|_{2}$, indicating the overall alignment stress. (2) Tension imbalance: $\mathrm{Var}(\vec{\tau}(p,t))$, measuring the anisotropy of tension distribution. (3) Tension skew: $\mathrm{Softmax}(\vec{\tau})$, yielding a probability-like attribution of hallucination directionality. Tracking $\vec{\tau}$ across diffusion steps reveals dynamic shifts in alignment priorities and exposes emerging risks of hallucination. As illustrated in Figure 4, certain prompts induce consistent dominance in specific $\tau_{i}$ components, leading to failure modes that are not random but axis-dependent. The ARC formulation thus transforms latent alignment pressures into a structured, actionable representation. In the next section, we leverage this vector to develop a control mechanism that dynamically mitigates hallucinations by counteracting emerging tensions in real time, without requiring retraining of the base diffusion model.
@@ 4 TensionModulator: ARC-Guided Hallucination Control Mechanism
While the ARC provides a real-time representation of cognitive tension during generation, it also enables a closed-loop correction framework. We introduce TensionModulator (TM-ARC), a lightweight, modular controller that applies dynamic perturbations to the generative trajectory in response to instantaneous tension signals. TM-ARC transforms ARC from a descriptive diagnostic vector into an actionable feedback mechanism, actively preventing directional drift toward the ideal manifold $\mathcal{M}_{\text{ideal}}$.
@@ 4.2 Modular Decomposition by Tension Axes
TM-ARC decomposes into three orthogonal sub-controllers, each aligned with a distinct cognitive alignment dimension: (1) SC-Gate ($F_{SC}$): Reactivates attention on prompt-critical entities when semantic tension $\tau_{SC}$ intensifies, counteracting semantic drift through modulation of decoder focus. (2) SA-Tuner ($F_{SA}$): Refines latent spatial encodings when structural tension $\tau_{SA}$ rises, preserving object placement and spatial logic via positional re-weighting. (3) KG-Augment ($F_{KG}$): Injects auxiliary factual priors or semantic anchors when knowledge tension $\tau_{KG}$ surges, reinforcing commonsense consistency. It operates in two complementary modes: (1) Static injection: KG-related prompt embeddings are prepended to the text encoder input to provide prior knowledge. (2) Dynamic modulation: Cross-attention layers are reweighted based on $\tau_{KG}$ to emphasize or suppress contextually relevant facts during generation. Each submodule operates as a soft perturbation function over the latent state $z_{t}$, guided solely by the corresponding ARC component. The modular structure ensures that each intervention remains interpretable and independently tunable.
@@ 4.3 ARC-Driven Feedback Dynamics
(1) $\lambda(\vec{\tau})=\sigma(\|\vec{\tau}\|_{2})$ is a tension-adaptive scaling function that increases modulation strength with overall tension magnitude; (2) $F_{i}(z_{t},\tau_{i})$ denotes the $i$-th correction operator, each modulated by its corresponding tension signal; (3) Each $F_{i}$ is differentiable and temporally localized, ensuring responsiveness without introducing global disruption. This formulation allows TM-ARC to function analogously to a physical damping system: minimal corrections under low tension, and strong directional feedback when misalignment intensifies.
@@ 4.4 Integration and Generalization
TM-ARC is designed as a plug-and-play augmentation for pretrained diffusion models. It operates entirely in the latent space and requires no additional supervision, training, or architecture modification. The modular design ensures compatibility across model backbones and preserves zero-shot generalizability. Because its intervention is conditioned solely on ARC: a model-internal, prompt-dependent signal, TM-ARC generalizes across unseen prompts and domains, as long as alignment tension can be monitored. By continuously sensing the evolving semantic, structural, and knowledge alignment tensions during generation, TM-ARC actively intervenes to guide the generative path back toward a cognitively balanced state. Thus, TM-ARC enables the first tension-aware feedback mechanism for hallucination mitigation in T2I diffusion, grounded in cognitive alignment dynamics rather than post hoc filtering or static constraints.
Table 1: Component-wise ablation of TM-ARC submodules. $F_{SC}$, $F_{SA}$, and $F_{KG}$ denote SC-Gate, SA-Tuner, and KG-Augment, respectively. Progressive activation shows complementary gains across CLIP, PickScore, ImageReward, and FID.
$F_{SC}$ $F_{SA}$ $F_{KG}$ CLIP Score $\uparrow$ PickScore $\uparrow$ ImageReward $\uparrow$ FID $\downarrow$
27.42 20.24 0.83 29.20
28.89 20.79 0.90 26.26
28.35 21.11 0.88 22.81
29.48 21.38 0.94 21.10
Table 2: Prompt-level ARC modulation improves generation faithfulness. ARC = $\tau_{SC}$, $\tau_{SA}$, $\tau_{KG}$]. Faithfulness improvement is observed after alignment tension correction.
Prompt (Hallucination-prone) ARC (Before) ARC (After) Faith $\uparrow$
A flying elephant in a business meeting [0.91, 0.27, 0.52] [0.48, 0.26, 0.33] +14.3
A red triangle with feathers walking upstairs [0.34, 0.86, 0.22] [0.29, 0.42, 0.21] +9.7
A cat reads a newspaper inside a microwave [0.49, 0.68, 0.77] [0.33, 0.41, 0.42] +12.1
A transparent piano floating above a desert war zone [0.73, 0.59, 0.71] [0.41, 0.34, 0.45] +13.5
A man balances Saturn on his fingertip while riding a dolphin [0.87, 0.39, 0.65] [0.52, 0.31, 0.37] +11.2
@@ 5 Experiments
5.1 3D ARC Enables Superior Unsupervised Clustering of Hallucination Types
We first assess whether ARC’s 3D representation enables meaningful unsupervised separation of hallucination types (Table 3). While prior analyses (e.g., CLIP or PCA embeddings) provide high-dimensional visual features, they lack interpretability and fail to cluster hallucinations into coherent categories. In contrast, the ARC vector $\bm{\tau}=[\tau_{SC},\tau_{SA},\tau_{KG}]$ is explicitly designed to reflect alignment tensions, offering a compact and cognitively grounded latent space.
Table 3: Unsupervised clustering performance ($k=3$) on DrawBench hallucination types. ARC-3D outperforms all baselines across ARI, NMI, and Accuracy.
Method ARI $\uparrow$ NMI $\uparrow$ Accuracy $\uparrow$
Random-3D (Noise) 0.09 0.14 0.37
PCA-50D (CLIP) 0.24 0.32 0.51
CLIP-3D 0.33 0.41 0.58
ARC-3D (Ours) 0.67 0.72 0.84
To evaluate this, we perform $k$-means clustering ($k=3$) on 300 generated images from the DrawBench benchmark Saharia et al. (2022), each labeled by majority vote from three annotators as exhibiting semantic, structural, or knowledge-level hallucination. Inter-annotator agreement is high ($\kappa=0.92$), and the class distribution is approximately balanced. ARC-3D achieves the highest scores across all metrics (ARI, NMI, Acc), despite using only 3 dimensions. We compare ARC-3D against several vision-only baselines: (1) Random-3D: Gaussian 512D noise projected to 3D. (2) PCA-50D: CLIP-Image features reduced to 50D via PCA. (3) CLIP-3D: CLIP-Image features reduced to 3D via PCA. (4) ARC-3D: Direct use of $\tau_{SC},\tau_{SA},\tau_{KG}$]. These results reinforce ARC’s effectiveness in structuring hallucination patterns along cognitively meaningful axes. Notably, even high-dimensional CLIP-based representations (PCA-50D) underperform relative to ARC’s low-dimensional encoding. This highlights that hallucinations are not only visually manifest, but are best understood through alignment tension decomposition. Implementation and labeling details are provided in Appendix A.
5.2 Component-wise Ablation Validates the Necessity of Multi-axis Tension Control
We conduct an ablation study to evaluate the contribution of each TM-ARC submodule in hallucination mitigation. As shown in Table 1, the baseline configuration (all modules disabled) achieves moderate performance. Activating $F_{SC}$ alone improves CLIP Score (+1.47) and ImageReward (+0.07), mitigating semantic drift. Adding $F_{SA}$ reduces FID by 3.45 points, improving spatial coherence. The full configuration ($F_{SC}+F_{SA}+F_{KG}$) yields the best performance, with $F_{KG}$ further improving ImageReward (+0.06) and FID (-1.71), ensuring factual consistency. This shows that $F_{SC}$ addresses semantic quality, $F_{SA}$ enhances structural alignment, and $F_{KG}$ improves factual consistency, confirming that hallucinations result from imbalanced tensions across semantic, structural, and knowledge dimensions.
@@ 5.3 Prompt-Level ARC Modulation Substantially Boosts Faithfulness
We assess ARC’s role in enabling targeted hallucination mitigation. As shown in Table 2, TM-ARC significantly reduces alignment risks across varied prompts. For example, “A flying elephant in a business meeting” shows high SC and KG tensions, which drop sharply after modulation, improving faithfulness by +14.3. Similarly, prompts with dominant SA tension (e.g., “A red triangle with feathers walking upstairs”) or multi-axial risk (e.g., “A transparent piano over a desert war zone”) also benefit from axis-specific attenuation. These results highlight ARC’s capacity to disentangle hallucination factors and TM-ARC’s ability to modulate them in a fine-grained, interpretable manner.
@@ 5.4 ARC Generalizes Across Backbones and Outperforms Strong Baselines
Table 4: Evaluation results on DrawBench and Pick-a-Pic datasets. Best results are bolded; second-best are underlined. ↑ indicates higher is better; ↓ indicates lower is better.
DrawBench Pick-a-Pic
Method CLIP Score↑ Pick Score↑ Image Reward↑ FID↓ CLIP Score↑ Pick Score↑ Image Reward↑ FID↓
UNet-based Models
SDXL
Vanilla Diffusion 27.18 20.66 0.54 31.07 27.34 20.11 0.64 29.93
Prompt-to-Prompt 28.33 20.85 0.58 28.43 29.88 20.41 0.70 29.75
Attend-and-Excite 28.67 21.14 0.65 24.35 28.86 20.77 0.83 25.33
Zigzag Diff 28.92 21.04 0.91 22.48 28.69 21.90 0.86 20.50
ARC (Ours) 29.26 21.60 0.85 20.02 29.21 22.19 0.90 21.47
SD1.5
Vanilla Diffusion 26.90 20.22 0.55 33.17 27.31 20.08 0.53 28.71
Prompt-to-Prompt 28.11 20.60 0.54 25.29 28.07 20.39 0.55 25.58
Attend-and-Excite 28.41 20.71 0.51 25.56 28.39 20.71 0.61 23.03
Zigzag Diff 28.51 20.63 0.62 23.34 28.59 21.02 0.54 19.99
ARC (Ours) 28.74 21.01 0.64 20.05 28.66 21.12 0.59 19.80
DiT-based Models
PixArt-sigma
Vanilla Diffusion 27.23 20.84 0.50 29.24 27.99 21.04 0.61 27.44
Prompt-to-Prompt 28.04 21.21 0.55 28.70 28.42 21.57 0.65 27.23
Attend-and-Excite 28.58 21.59 0.69 28.90 29.11 21.69 0.79 25.98
Zigzag Diff 29.10 21.25 0.75 24.63 29.44 21.21 0.81 20.97
ARC (Ours) 28.99 21.99 0.79 19.84 29.25 21.72 0.83 17.73
Hunyuan-DiT
Vanilla Diffusion 27.59 21.14 0.81 30.22 27.80 21.69 0.85 29.88
Prompt-to-Prompt 28.94 21.71 0.84 23.43 28.31 22.16 0.92 24.09
Attend-and-Excite 28.79 22.54 0.89 21.95 28.71 22.68 0.95 21.44
Zigzag Diff 29.09 21.61 0.79 19.20 28.94 20.83 0.95 19.25
ARC (Ours) 29.94 22.59 0.92 18.17 29.73 21.54 0.98 18.88